# Equivariant scaling asymptotics for Poisson and Szegő kernels on Grauert tube boundaries

### v1

@@ 1.2.2 Scaling asymptotics for $P^{\tau}_{\chi,\nu,\lambda}$
    the leading order terms of the asymptotic expansions in Theorems 25, 15 and 28 are multiplied by $(\lambda/\tau)^{-(d-1)/2}$;
@@
$\sim$ $\frac{1}{\sqrt{2\,\pi}}\cdot\left(\frac{1}{2\,\pi}\right)^{d-1}\cdot\left(\frac{\lambda}{\tau}\right)^{\frac{d-1}{2}}\cdot e^{\frac{1}{\tau}\,\left[\imath\,\sqrt{\lambda}\,(\theta_{1}-\theta_{2})+\Psi_{M^{-1}}(\mathbf{v}_{1},\mathbf{v}_{2})\right]}\cdot e^{\imath\,\tilde{\theta}^{\tau}_{t_{1}}(x)}\,$
$\cdot\left[e^{-\imath\,\lambda\,t_{1}}\cdot\chi(t_{1})+\sum_{k\geq 1}\,\lambda^{-k/2}\,\widetilde{F}_{k}\left(x_{1},x_{2};\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2}\right)\right],$
for certain polynomials $\widetilde{F}_{k}(x_{1},x_{2};\cdot)$ in the rescaled variables of degree $\leq 3\,k$ and parity $k$. The expansion in Theorem 1.9 of [CR2] follows by rescaling according to Heisenberg type (as in the Introduction of [P-2024]).
@@ 1.3 Unrescaled Asymptotics
    $\cdot\left[e^{\imath\,\theta^{\tau}_{t_{1}}(x_{1})}\cdot\mathcal{F}_{\chi,\lambda}(x_{1},x_{2})\cdot\mathcal{B}_{\nu}(x_{1},x_{2})_{l}+\sum_{s\geq 1}\,\lambda^{-s}\,L_{\nu,l,s}\left(x_{1},x_{2};\theta_{1},\mathbf{v}_{1}^{t},\theta_{2},\mathbf{v}_{2}^{t}\right)\right],$
@@
    $\cdot\left[e^{\imath\,\theta^{\tau}_{t_{1}}(x_{1})}\cdot\mathcal{F}_{\chi,\lambda}(x_{1},x_{2})\cdot\mathcal{B}_{\nu}(x_{1},x_{2})_{l}+\sum_{s\geq 1}\,\lambda^{-s}\,K_{\nu,l,s}\left(x_{1},x_{2};\theta_{1},\theta_{2},\mathbf{V}\right)\right],$
@@ 1.4.1 An equivariant global Weyl law for $\mathfrak{D}^{\tau}_{\sqrt{\rho}}$
$\mathcal{W}_{\nu}^{\tau}(\lambda)$ $=$ $\frac{1}{2^{d_{G}}}\cdot\frac{\tau}{d-d_{G}}\cdot\left(\frac{\lambda}{2\,\pi\,\tau}\right)^{d-d_{G}}\cdot\dim(\nu)^{2}\cdot\mathrm{vol}(Z^{\tau}/G)$
@@ 1.4.3 Operator norm estimates for $\Pi^{\tau}_{\chi,\nu,\lambda}$
This result is the analogue of similar estimates proved in the line bundle setting by Shiffman and Zelditch in [SZ03], and the proof follows the same general lines, adapted to the Grauert tube context. More precisely, the argument given by Chang and Rabinowitz for (29) is based on the off-graph scaling asymptotics of $\Pi^{\tau}_{\chi,\lambda}$, paired with the Shur-Young inequality [So].
@@
$\sim$ $\frac{\sqrt{2\,\pi}\cdot\pi^{d_{G}}}{(2\,\pi)^{d}}\,\left(\frac{\lambda}{\tau}\right)^{\frac{d-1}{2}-d_{G}}\,\cdot\dim(\nu)^{2}\cdot\mathrm{vol}(Z^{\tau}/G)$
@@
There is a constanct $C_{\nu}^{\tau}>0$ such that for all $j\gg 0$
@@
$\mathcal{P}_{\nu}^{\tau}(\lambda)$ $\sim$ $\frac{\pi^{d_{G}}}{(2\,\pi)^{d}}\,\left(\frac{\lambda}{\tau}\right)^{\frac{d-1}{2}-d_{G}}\,\cdot\dim(\nu)^{2}\cdot\mathrm{vol}(Z^{\tau}/G)$
@@ 2 Preliminaries
We shall identify the tangent and cotangent bundles of $M$, $TM$ and $T^{\vee}M$, by means of $\kappa$.
@@ 2.1 The action on the Grauert tubes
In Section 1, the Grauert tubes $\tilde{M}^{\tau}$, $\tau\in(0,\tau_{0})$, have been described as strictly pseudoconvex domains in $\tilde{M}$ intrisically associated to $\kappa$. An alternative characterization of $\tilde{M}^{\tau}$ is in terms of disk bundles $T^{\tau}M\subset T^{\vee}M$ of norm $\tau$ with respect to $\kappa$, where $T^{\vee}M$ has been endowed in a neighbourhood of the zero section $M_{0}$ with a complex structure intrinsically associated to $\kappa$ ([LS], [Lem], [Sz], [GS-1991], [GS-1992]). This complex structure was called adapted in [LS], and will be denoted $J_{ad}$; it is uniquely determined by the condition that the Riemannian and Monge-Ampère foliations coincide [LS].
@@
Since $\rho$ is $G$-invariant, $[\xi_{\tilde{M}^{\tau_{0}}},\upsilon_{\sqrt{\rho}}]=0$, hence $\Phi$ is constant along the geodesic flow. Recalling (7) and (19), we conclude the following.
@@ 2.2 $\Pi^{\tau}$ in NHLC’s on $X^{\tau}$
    $s^{\tau}$ is a classical symbol of type:
@@
Suppose $x\in X^{\tau}$ and fix a NHLC’s on $X^{\tau}$ centered at $x$ and defined on an open neighourhood $U^{\tau}\subseteq X^{\tau}$ of $x$. Then, if $x^{\prime},x^{\prime\prime}\in U^{\tau}$ are of the form $x^{\prime}=x+(\theta,z^{\prime})$, $x^{\prime\prime}=x+(\eta,z^{\prime\prime})$, we have:
@@
Moreover, in the same chart, the principal term of the symbol along the diagonal takes the form
@@ 2.3 The submanifold $Z^{\tau}$
Our analysis rests on the assumption that $G$ acts locally freely on $Z^{\tau}$ in (19). This hypothesis is certainly satisfied if $\mu$ itself is locally free - for instance if $M$ is the total space of a principal $G$-bundle. This is not the only circumstance. Here we give two sufficient conditions, whose proof is left to the reader.
@@
If $G$ acts locally freely on $Z\setminus M$, the latter is a submanifold of $\tilde{M}$ of codimension $d_{G}$. Given that $Z$ is obviously transverse to $X^{\tau}$, we conclude the following.
@@ 2.4 Intersections of $G$-orbits and $\mathbb{R}$-orbits
    There exists $R,\,r>0$ such that if $(g,t)$ belongs to an $R$-neighbourhood of $(e_{G},0)$ then
@@ 2.5 The action in Heisenberg local coordinates
As is the Introduction (recall the discussion surrounding (21)), NHLC’s on $X^{\tau}$ centered at $x$ will be denoted in additive notation: $y(\theta,\mathbf{v})=x+(\theta,\mathbf{v})$. Given $x\in X^{\tau}$, we can find an open neighbourhood $U^{\tau}\subseteq X^{\tau}$ of $x$ and a smoothly varying family of systems of NHLC’s $y_{x^{\prime}}(\theta,\mathbf{v})=x^{\prime}+(\theta,\mathbf{v})$ centered at $x^{\prime}\in X^{\tau}$. More explicitly, for some $\delta>0$ we have a smooth map $\Upsilon:U^{\tau}\times(-\delta,\delta)\times B_{2d-2}(\mathbf{0},\delta)\rightarrow X^{\tau}$ such that for every $x^{\prime}\in U^{\tau}$ the partial function $\Upsilon(x^{\prime},\cdot,\cdot)$ is a system of NHLC’s centered at $x^{\prime}$. Hence
@@
$-\imath\,\beta_{2}(\theta_{j},\mathbf{u}_{j})-\frac{1}{2}\,\|\mathbf{u}_{1}-\mathbf{u}_{2}\|^{2}=\psi_{2}(\mathbf{u}_{1},\mathbf{u}_{2})=-\imath\,\omega_{x}(\mathbf{u}_{1},\mathbf{u}_{2})-\frac{1}{2}\,\|\mathbf{u}_{1}-\mathbf{u}_{2}\|^{2}.$
@@ 3.1 Preamble
    $\tilde{\Pi}^{\tau}_{t}(x,y):=\left(\Gamma^{\tau}_{-t}\times\Gamma^{\tau}_{-t}\right)^{*}\left(\Pi^{\tau}(x,y)\right)=\Pi^{\tau}\left(\Gamma^{\tau}_{-t}(x),\Gamma^{\tau}_{-t}(y)\right).$
@@ 3.2 The proof
Localizing the computation near some $\left(\tilde{g},\tilde{t}\right)$, we shall set $\tilde{x}:=\mu^{\tau}_{\tilde{g}^{-1}}(x_{1})$ (so that there is a naturally induced system of NHLC’s centered at $\tilde{x}$) and
@@
Let us fix a system of NHLC’s at $\Gamma^{\tau}_{-\tilde{t}}(\tilde{x})$. Furthermore, arguing as in Lemma 67, for a suitable symplectic matrix $M_{\tilde{t}}$, we obtain
@@
The variable $\theta$ being also compactly supported, and $v$ ranges in a certain neighbourhood of $1/\tau$; therefore, upon choosing $\epsilon$ and $\delta$ suitably small, we conclude that the contribution to the asymptotics of the locus where $0<u\ll 1/\tau$ or $u\gg 1/\tau$ is also negligible. In other words, the asymptotics of (3.2) are unaltered, if the integrand is multiplied by a cut-off function of the form $f_{2}(u)$, where $f_{2}\in\mathcal{C}^{\infty}_{c}(\mathbb{R}_{+})$ is identically equal to $1$ on an appropriate neighbourhood of $1/\tau$.
@@
Since $u$ is now bounded away from zero, we conclude by iteratively integrating by parts in $g$ that the contribution to the asymptotics of the locus where $\|\Phi(x_{1})\|\leq C\,\lambda^{\epsilon^{\prime}-\frac{1}{2}}$ is negligible. Given that by assumption 00 is a regular value of $\Phi$, we conclude that
$\Pi^{\tau}_{\chi,\nu,\lambda}(x_{1},x_{2})=O\left(\lambda^{\epsilon^{\prime}-1/2}\right),$
@@ 4.1 General arguments for the scaling asymptotics of $\Pi^{\tau}_{\chi,\nu,\lambda}$
Let $\gamma^{\mathfrak{g}}:\mathfrak{g}\rightarrow\mathbb{R}$ denote a bump function supported in a small neighbourhood of $0\in\mathfrak{g}$, and identically equal to $1$ sufficiently close to 00. Similarly, let $\gamma^{\mathbb{R}}:\mathbb{R}\rightarrow\mathbb{R}$ denote a compactly supported bump function compactly supported on a small neighbourhood of $0\in\mathbb{R}$, and identically equal to $1$ sufficiently close to 00. We conclude that the asymptotics of $\Pi^{\tau}_{\chi,\nu,\lambda}(x_{1,\lambda},x_{2,\lambda})$ are unaltered, if the integrand in (4.1) is further multiplied by the expression
@@
where $(t,\xi)$ now ranges in a neighbourhood of the origin in ×$\mathbb{R}\times\mathfrak{g}$. The $l$-th summand in (81) now takes the form
@@
so that integration in $(\xi,t)$ is now over a ball centered at the origin and radius $O\left(\lambda^{\epsilon^{\prime}}\right)$ in $\mathfrak{g}\times\mathbb{R}$. Similarly, since prior to rescaling $(\theta,\mathbf{v})$ range in a shrinking ball centered at the origin of radius $O\left(\lambda^{\epsilon^{\prime}-1/2}\right)$, the rescaled variables will range in an expanding ball of radius $O\left(\lambda^{\epsilon^{\prime}}\right)$ in ×$\mathbb{R}\times\mathbb{R}^{2d-2}$.
@@
::=:= $\Psi\left(x_{1,\lambda},x_{2,\lambda};e^{\xi/\sqrt{\lambda}}\,\kappa_{l}\,h_{1},t_{1}+\frac{t}{\sqrt{\lambda}},y_{\lambda}(\theta,\mathbf{u}),u,v\right),$
@@
::=:= $\Xi_{\nu}\left(h_{1}^{-1}\,\kappa_{l}^{-1}\,e^{-\xi/\sqrt{\lambda}}\right)\,\chi\left(t_{1}+\frac{t}{\sqrt{\lambda}}\right)\cdot\mathcal{V}\left(\frac{\theta}{\sqrt{\lambda}},\frac{\mathbf{u}}{\sqrt{\lambda}}\right)$
@@
::=:= $u\,\left[\imath\,\omega_{0}\big{(}\xi_{X^{\tau}}(x_{1}),\mathbf{v}_{1}\big{)}-\frac{1}{4\,\tau^{2}}\,(\theta_{1}-\theta)^{2}+\psi_{2}\left(\mathbf{v}^{(l)}_{1}-\xi_{X^{\tau}}(x_{1})^{(l)},\mathbf{u}\right)\right]$
@@
::=:= $e^{\mathcal{S}(\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2};t,v,\theta,u,\mathbf{u},\xi)_{l}}\,e^{\lambda\,R_{3}\left(\frac{\bullet}{\sqrt{\lambda}}\right)}\cdot\mathcal{A}_{\lambda}(\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2},t,\xi,\theta,\mathbf{u},u,v)_{l}.$
@@
::=:= $\int_{-\infty}^{+\infty}\,\mathrm{d}t\,\int_{-\infty}^{+\infty}\,\mathrm{d}\theta\,\int_{0}^{+\infty}\,\mathrm{d}u\,\int_{0}^{+\infty}\,\mathrm{d}v$
@@
Leaving a corresponding cut-off function in $(\theta,t)$ implicit, we may now study the asymptotics of $I_{\lambda}(x_{j},\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2},\xi,\mathbf{u})_{l}$ in (4.1) using the stationary phase Lemma. The phase (4.1) has already been considered in [Grauert-9-2023] (where it is denoted $\Upsilon^{\tau}$). By Lemma 64 of the same paper, (94) has a unique stationary point $P_{s}=(t_{s},v_{s},\theta_{s},u_{s})$, given by
@@
furthermore, the Hessian determinant and signature at the critical point are, respectively, $\tau^{2}$ and 00. Arguing as in §4 of [Grauert] (in particular, as in the derivation of (147) in loc. cit.) we then obtain an asymptotic expansion for $I_{\lambda}(x_{j},\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2},\xi,\mathbf{u})_{l}$ in (4.1) of the following form:
@@
::=:= $\frac{1}{\tau}\,\left[\imath\,\omega_{0}\big{(}\xi_{X^{\tau}}(x_{1}),\mathbf{v}_{1}\big{)}+\psi_{2}\left(\mathbf{v}^{(l)}_{1}-\xi_{X^{\tau}}(x_{1})^{(l)},\mathbf{u}\right)+\psi_{2}\big{(}B\mathbf{u},\mathbf{v}_{2}\big{)}\right],$

### v3

@@ 1.1.4 Scaling asymptotics for $P^{\tau}_{\chi,\nu,\lambda}$
    the leading order terms of the asymptotic expansions in Theorems 25, 15 and 28 are multiplied by $(\lambda/\pi\,\tau)^{-(d-1)/2}$;
@@
$\sim$ $\frac{e^{-\imath\,\lambda\,t_{1}}}{\sqrt{2\,\pi}}\cdot\left(\frac{1}{2}\right)^{d-1}\cdot\left(\frac{\lambda}{\tau\,\pi}\right)^{\frac{d-1}{2}}\cdot e^{\frac{1}{\tau}\,\left[\imath\,\sqrt{\lambda}\,(\theta_{1}-\theta_{2})+\Psi_{M^{-1}}(\mathbf{v}_{1},\mathbf{v}_{2})\right]}\cdot e^{\imath\,\tilde{\theta}^{\tau}_{t_{1}}(x)}\,$
$\cdot\left[\chi(t_{1})+\sum_{k\geq 1}\,\lambda^{-k/2}\,\widetilde{F}_{k}\left(x_{1},x_{2};\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2}\right)\right],$
for certain polynomials $\widetilde{F}_{k}(x_{1},x_{2};\cdot)$ in the rescaled variables of degree $\leq 3\,k$ and parity $k$. In the yet more particular near-diagonal case, the previous expansion corrects the leading order factor appearing in the expansion in Theorem 7 of [P-2024] by a power of $\pi^{(d-1)/2}$; this is related to the symbolic computation in §5 below (see Lemma 81).
The original expansion in Theorem 1.9 of [CR2] follows by rescaling according to Heisenberg type (as in the Introduction of [P-2024]).
@@ 1.2 Unrescaled Asymptotics
    $\cdot\left[e^{\imath\,\theta^{\tau}_{t_{1}}(x_{1})}\cdot\mathcal{F}_{\chi}(x_{1},x_{2})\cdot\mathcal{B}_{\nu}(x_{1},x_{2})_{l}+\sum_{s\geq 1}\,\lambda^{-s}\,L_{\nu,l,s}\left(x_{1},x_{2};\theta_{1},\mathbf{v}_{1}^{t},\theta_{2},\mathbf{v}_{2}^{t}\right)\right],$
@@
    $\cdot\left[e^{\imath\,\theta^{\tau}_{t_{1}}(x_{1})}\cdot\mathcal{F}_{\chi}(x_{1},x_{2})\cdot\mathcal{B}_{\nu}(x_{1},x_{2})_{l}+\sum_{s\geq 1}\,\lambda^{-s}\,K_{\nu,l,s}\left(x_{1},x_{2};\theta_{1},\theta_{2},\mathbf{V}\right)\right],$
@@ 1.3.1 An equivariant global Weyl law for $\mathfrak{D}^{\tau}_{\sqrt{\rho}}$
$\mathcal{W}_{\nu}^{\tau}(\lambda)$ $=$ $\frac{1}{2^{d_{G}/2}}\cdot\frac{\tau}{d-d_{G}}\cdot\left(\frac{\lambda}{2\,\pi\,\tau}\right)^{d-d_{G}}\cdot\dim(\nu)^{2}\cdot\mathrm{vol}(Z^{\tau}/G)$
@@ 1.3.3 Operator norm estimates for $\Pi^{\tau}_{\chi,\nu,\lambda}$
This result is the analogue of similar estimates proved in the line bundle setting by Shiffman and Zelditch in [SZ03], and the proof follows the same general lines, adapted to the Grauert tube context (for antecedents in the real domain, see [So-1988] and [So-2017]). More precisely, the argument given by Chang and Rabinowitz for (29) is based on the off-graph scaling asymptotics of $\Pi^{\tau}_{\chi,\lambda}$, paired with the Shur-Young inequality [So-2017].
@@
$\sim$ $\frac{1}{\sqrt{2^{d+d_{G}}\,\pi}}\,\left(\frac{\lambda}{2\,\pi\,\tau}\right)^{\frac{d-1}{2}-d_{G}}\,\cdot\dim(\nu)^{2}\cdot\mathrm{vol}(Z^{\tau}/G)$
@@
There is a constant $C_{\nu}^{\tau}>0$ such that for all $j\gg 0$
@@
$\mathcal{P}_{\nu}^{\tau}(\lambda)$ $\sim$ $\frac{1}{\sqrt{2^{d+1+d_{G}}}\,\pi}\,\left(\frac{\lambda}{2\,\pi\,\tau}\right)^{\frac{d-1}{2}-d_{G}}\,\cdot\dim(\nu)^{2}\cdot\mathrm{vol}(Z^{\tau}/G)$
@@ 2 Preliminaries
We shall identify the tangent and cotangent bundles of $M$, TMTM and $T^{\vee}M$, by means of $\kappa$.
@@ 2.1 The action on the Grauert tubes
In Section 1, the Grauert tubes $\tilde{M}^{\tau}$, $\tau\in(0,\tau_{0})$, have been described as strictly pseudoconvex domains in $\tilde{M}$ intrisically associated to $\kappa$. An alternative characterization of $\tilde{M}^{\tau}$ is in terms of disk bundles $T^{\tau}M\subset T^{\vee}M$ of norm $\tau$ with respect to $\kappa$, where $T^{\vee}M$ has been endowed in a neighbourhood of the zero section $M_{0}$ with a complex structure intrinsically associated to $\kappa$ ([LS], [Lem], [Sz], [GStz-1991], [GStz-1992]). This complex structure was called adapted in [LS], and will be denoted $J_{ad}$; it is uniquely determined by the condition that the Riemannian and Monge-Ampère foliations coincide [LS].
@@
Since $\rho$ is $G$-invariant, $\xi_{\tilde{M}^{\tau_{0}}},\upsilon_{\sqrt{\rho}}]=0$, hence $\Phi$ is constant along the geodesic flow. Recalling (7) and (19), we conclude the following.
@@ 2.2 $\Pi^{\tau}$ in NHLC’s on $X^{\tau}$
    $s^{\tau}$ is a classical symbol of the form:
@@
[…]
@@
Moreover, in the same chart, the principal term of the symbol satisfies
@@ 2.3 The submanifold $Z^{\tau}$
Our analysis rests on the assumption that $G$ acts locally freely on $Z^{\tau}$ in (19). Thus Assumption 4 is certainly satisfied if $\mu$ itself is locally free and $Z^{\tau}\neq\emptyset$, for instance if $M$ is the total space of a principal $G$-bundle over a non-trivial base, $p:M\rightarrow N$ with $\dim(N)>0$. In this case, under the map $E$ in §2.1, $Z$ is the image in $\tilde{M}$ of the orthocomplement of the vertical tangent bundle $\mathrm{Ver}(p)\subseteq TM$ (that is, of the horizontal tangent bundle $\mathrm{Hor}(p)\subseteq TM$) of $M$ as a principal $G$-bundle (more precisely, $Z=E\big{(}\mathrm{Hor}(p)\big{)}\cap T^{\prime}M$). Then the quotient $Z/G$ can be identified with a tubular neighbourhood of $N$ in $TN$, and $Z^{\tau}/G$ with the bundle of tangent spheres of radius $\tau$ over $N$.
The case of a principal $G$-bundle is clearly not the only circumstance where the hypothesis is statisfied; for example, if $G$ acts in a Hamiltonian manner on a compact real-analytic symplectic manifold $(R,\omega_{R})$, and $0\in\mathfrak{g}^{\vee}$ is a regular value of the moment map $\Phi_{R}:R\rightarrow\mathfrak{g}^{\vee}$, then $G$ acts locally freely on $M:=\Phi_{R}^{-1}(0)$, but in most cases the action is not free.
Furthermore, it can happen that Assumption 4 is satisfied even if $\mu$ is not locally free. Here we give two more sufficient conditions, whose proof is left to the reader.
@@
If $G$ acts locally freely on $Z\setminus M$, the latter is a submanifold of $\tilde{M}$ of codimension $d_{G}$ (Remark 5). Given that $Z$ is obviously transverse to $X^{\tau}$, we conclude the following.
@@ 2.4 Intersections of $G$-orbits and $\mathbb{R}$-orbits
    There exists R,r>0R,\,r>0 such that if $(g,t)$ belongs to an $R$-neighbourhood of $(e_{G},0)$ then
@@ 2.5 The action in Heisenberg local coordinates
As is the Introduction (recall the discussion surrounding (21)), NHLC’s on $X^{\tau}$ centered at $x$ will be denoted in additive notation: $y(\theta,\mathbf{v})=x+(\theta,\mathbf{v})$. Given $x\in X^{\tau}$, we can find an open neighbourhood $U^{\tau}\subseteq X^{\tau}$ of $x$ and a smoothly varying family of systems of NHLC’s $y_{x^{\prime}}(\theta,\mathbf{v})=x^{\prime}+(\theta,\mathbf{v})$ centered at $x^{\prime}\in U^{\tau}$. More explicitly, for some $\delta>0$ we have a smooth map $\Upsilon:U^{\tau}\times(-\delta,\delta)\times B_{2d-2}(\mathbf{0},\delta)\rightarrow X^{\tau}$ such that for every $x^{\prime}\in U^{\tau}$ the partial function $\Upsilon(x^{\prime},\cdot,\cdot)$ is a system of NHLC’s centered at $x^{\prime}$. Hence
@@
$\imath\,\beta_{2}(\theta_{j},\mathbf{u}_{j})-\frac{1}{2}\,\|\mathbf{u}_{1}-\mathbf{u}_{2}\|^{2}=\psi_{2}(\mathbf{u}_{1},\mathbf{u}_{2})=-\imath\,\omega_{x}(\mathbf{u}_{1},\mathbf{u}_{2})-\frac{1}{2}\,\|\mathbf{u}_{1}-\mathbf{u}_{2}\|^{2}.$
@@ 3.1 Preamble
    $\tilde{\Pi}^{\tau}_{t}(x,y):=\left(\Gamma^{\tau}_{-t}\times\Gamma^{\tau}_{-t}\right)^{*}\left(\Pi^{\tau}\right)(x,y)=\Pi^{\tau}\left(\Gamma^{\tau}_{-t}(x),\Gamma^{\tau}_{-t}(y)\right).$
@@ 3.2 The proof
[…]
@@
[…]
@@
The variable $\theta$ is also compactly supported, and $v$ ranges in a certain neighbourhood of $1/\tau$; therefore, upon choosing $\epsilon$ and $\delta$ suitably small, we conclude that the contribution to the asymptotics of the locus where $0<u\ll 1/\tau$ or $u\gg 1/\tau$ is also negligible. In other words, the asymptotics of (3.2) are unaltered, if the integrand is multiplied by a cut-off function of the form $f_{2}(u)$, where $f_{2}\in\mathcal{C}^{\infty}_{c}(\mathbb{R}_{+})$ is identically equal to $1$ on an appropriate neighbourhood of $1/\tau$.
@@
Since $u$ is now bounded away from zero, we conclude by iteratively integrating by parts in $g$ that the contribution to the asymptotics of the locus where $\|\Phi(x_{1})\|\geq C\,\lambda^{\epsilon^{\prime}-\frac{1}{2}}$ is negligible. Given that by assumption 0 is a regular value of $\Phi$, we conclude that
$\Pi^{\tau}_{\chi,\nu,\lambda}(x_{1},x_{2})=O\left(\lambda^{-\infty}\right),$
@@ 4.1 General arguments for the scaling asymptotics of $\Pi^{\tau}_{\chi,\nu,\lambda}$
Let $\gamma^{\mathfrak{g}}:\mathfrak{g}\rightarrow\mathbb{R}$ denote a bump function supported in a small neighbourhood of $0\in\mathfrak{g}$, and identically equal to $1$ sufficiently close to 0. Similarly, let $\gamma^{\mathbb{R}}:\mathbb{R}\rightarrow\mathbb{R}$ denote a compactly supported bump function compactly supported on a small neighbourhood of $0\in\mathbb{R}$, and identically equal to $1$ sufficiently close to 0. We conclude that the asymptotics of $\Pi^{\tau}_{\chi,\nu,\lambda}(x_{1,\lambda},x_{2,\lambda})$ are unaltered, if the integrand in (4.1) is further multiplied by the expression
@@
where $(t,\xi)$ now ranges in a neighbourhood of the origin in $\mathbb{R}\times\mathfrak{g}$. The $l$-th summand in (81) now takes the form
@@
so that integration in $(\xi,t)$ is now over a ball centered at the origin and radius $O\left(\lambda^{\epsilon^{\prime}}\right)$ in $\mathfrak{g}\times\mathbb{R}$. Similarly, since prior to rescaling $(\theta,\mathbf{v})$ range in a shrinking ball centered at the origin of radius $O\left(\lambda^{\epsilon^{\prime}-1/2}\right)$, the rescaled variables will range in an expanding ball of radius $O\left(\lambda^{\epsilon^{\prime}}\right)$ in $\mathbb{R}\times\mathbb{R}^{2d-2}$.
@@
$:=$ $\Psi\left(x_{1,\lambda},x_{2,\lambda};e^{\xi/\sqrt{\lambda}}\,\kappa_{l}\,h_{1},t_{1}+\frac{t}{\sqrt{\lambda}},y_{\lambda}(\theta,\mathbf{u}),u,v\right),$
@@
$:=$ $\Xi_{\nu}\left(h_{1}^{-1}\,\kappa_{l}^{-1}\,e^{-\xi/\sqrt{\lambda}}\right)\,\chi\left(t_{1}+\frac{t}{\sqrt{\lambda}}\right)\cdot\mathcal{V}\left(\frac{\theta}{\sqrt{\lambda}},\frac{\mathbf{u}}{\sqrt{\lambda}}\right)$
@@
$:=$ $u\,\left[\imath\,\omega_{0}\big{(}\xi_{X^{\tau}}(x_{1}),\mathbf{v}_{1}\big{)}-\frac{1}{4\,\tau^{2}}\,(\theta_{1}-\theta)^{2}+\psi_{2}\left(\mathbf{v}^{(l)}_{1}-\xi_{X^{\tau}}(x_{1})^{(l)},\mathbf{u}\right)\right]$
@@
$:=$ $e^{\mathcal{S}(\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2};t,v,\theta,u,\mathbf{u},\xi)_{l}}\,e^{\lambda\,R_{3}\left(\frac{\bullet}{\sqrt{\lambda}}\right)}\cdot\mathcal{A}_{\lambda}(\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2},t,\xi,\theta,\mathbf{u},u,v)_{l}.$
@@
$:=$ $\int_{-\infty}^{+\infty}\,\mathrm{d}t\,\int_{-\infty}^{+\infty}\,\mathrm{d}\theta\,\int_{0}^{+\infty}\,\mathrm{d}u\,\int_{0}^{+\infty}\,\mathrm{d}v$
@@
Leaving a corresponding cut-off function in $(\theta,t)$ implicit, we may now study the asymptotics of $I_{\lambda}(x_{j},\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2},\xi,\mathbf{u})_{l}$ in (4.1) using the stationary phase Lemma. The phase (4.1) has already been considered in [P-2024] (where it is denoted $\Upsilon^{\tau}$). By Lemma 64 of the same paper, (94) has a unique stationary point $P_{s}=(t_{s},v_{s},\theta_{s},u_{s})$, given by
@@
furthermore, the Hessian determinant and signature at the critical point are, respectively, $\tau^{2}$ and 0. Arguing as in §4 of [P-2024] (in particular, as in the derivation of (147) in loc. cit.) we then obtain an asymptotic expansion for $I_{\lambda}(x_{j},\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2},\xi,\mathbf{u})_{l}$ in (4.1) of the following form:
@@
$:=$ $\frac{1}{\tau}\,\left[\imath\,\omega_{0}\big{(}\xi_{X^{\tau}}(x_{1}),\mathbf{v}_{1}\big{)}+\psi_{2}\left(\mathbf{v}^{(l)}_{1}-\xi_{X^{\tau}}(x_{1})^{(l)},\mathbf{u}\right)+\psi_{2}\big{(}B\mathbf{u},\mathbf{v}_{2}\big{)}\right],$